A potter’s wheel moves from rest to an angular speed of 0.10 rev/s in 30.9 s.
Assuming constant angular acceleration,
what is its angular acceleration in rad/s2?
Answer in units of rad/s2.
1
Expert's answer
2019-12-19T11:36:46-0500

The  angular acceleration:


"\alpha=\frac{\omega-0}{t}"

"\alpha=\frac{0.10}{30.9}=0.0032\frac{rad}{s^2}"